SpaceRef

SpaceRef


NASA ISS On-Orbit Status 22 November 2008

Status Report From: NASA HQ
Posted: Saturday, November 22, 2008

image

All ISS systems continue to function nominally, except those noted previously or below. Saturday – Flight Day 9 (FD9) of STS-126/ULF-2. ISS crew work cycle today: Wake 8:55am EST; sleep 12:25am; Shuttle crew: 12:55am (both until 8:55am tomorrow).

Mission ULF-2’s EVA-3 is underway, having begun at 1:01pm EST (44 min ahead of schedule). EVA-3 is being performed by Heidemarie Stefanyshyn-Piper (EV1) & Steve Bowen (EV2). [EV1 & EV2 began their “campout” (nachalo desaturatsiy = desaturation start) last night at ~11:20pm in the U.S. Airlock (A/L) with hatch closure and depressurization of the CL (Crewlock) from 14.7 to 10.2 psi, followed by mask prebreathe at ~11:20pm-12:25am. This morning, following the usual hygiene break/with mask prebreathe for the two spacewalkers at ~9:30am-10:40am after spending the night on 10.2 psi, the A/L hatch was closed again for EVA preps in 10.2 psi, followed by EMU (Extravehicular Mobility Unit) purge and prebreathe in the EMUs, assisted by CDR Ferguson. Afterwards, with CL depressurization and EV1/EV2 egress, EVA-3 began at 1:01pm, with MS4 Kimbrough providing IVA (Intravehicular Activity) assist.]

EVA-3 is expected to last about 7h (i.e., until ~8:01pm). Spacewalk objectives at the Stbd SARJ (Solar Alpha Rotary Joint) are –
For EV1/Piper:

  • Clean & lube under covers 15 & 16; install TBA (Trundle Bearing Assembly)-1 plus covers 15 & 16;
  • Remove, clean & lube under covers 13, 14, 11, 12, 17, 18; remove, clean, lube under TBAs-1,-7,-12,-3;
  • Install TBAs-1, -7, -12, -3; install covers 13, 14, 11, 12, 17, 18.

For EV2/Bowen:

  • Remove, clean & lube under covers 21, 22, 1(do not R&R TBA-6);
  • Remove, clean & lube under covers 2, 3, 19, 20; remove, clean & lube under TBA-7,-4;
  • Clean TBA-4 DLA (Drive Lock Assembly);
  • Install TBAs-7,-4; install covers 2, 3, 19, 20 (do not R&R TBA-5).

A NOAX (Non-Oxide Adhesive Experimental) caulk applicator from the Shuttle, using Braycote-601 lubricant for the trundle bearings, was approved for use on EVA-3 & EVA-4. [For EVA-3, it was left in the A/L for use if needed. For EVA-4, it will be modified to reduce Braycote flow from its nozzle.]

After the spacewalkers’ return on board tonight, post-EVA activities by MS3 Stefanyshyn-Piper, MS2 Bowen, FE-2 Chamitoff & CDR Ferguson in the A/L will consist of --

  • Recharging the EMU/spacesuits with water from PWR (Payload Water Reservoir),
  • Reconnecting the LTAs (Lower Torso Assemblies) to the EMUs
  • Initiating METOX (Metal Oxide) canister regeneration,
  • Capping the UIA (Umbilical Interface Assembly,
  • Initiating battery charging in the A/L BSA (Battery Stowage Assembly),
  • Taking photographs of the EMU gloves for downlink and inspection.

Before breakfast, FE-2-18 Sandra Magnus completed her first session with the biomed experiment INTEGRATED IMMUNE (Validating Procedures for Monitoring Crew member Immune Function), collecting a wet saliva sample. [IMMUNE protocol requires the collection to occur first thing post-sleep, before eating, drinking and brushing teeth, and all samples are stored at ambient temperature. Along with NUTRITION (Nutritional Status Assessment), INTEGRATED IMMUNE samples & analyzes participant’s blood, urine, and saliva before, during and after flight for changes related to functions like bone metabolism, oxidative damage and immune function to develop and validate an immune monitoring strategy consistent with operational flight requirements and constraints. The strategy uses both long and short duration crewmembers as study subjects. The saliva is collected in two forms, dry and liquid. The dry samples are collected at intervals during the collection day using a specialized book that contains filter paper. The liquid saliva collections require that the crewmember soak a piece of cotton inside their mouth and place it in a salivette bag; there are four of the liquid collections during docked operations.]

After having successfully deinstalled the K2-VKA-01 container of the KURS-P automated rendezvous radar system in the FGB, FE-1 Lonchakov is currently mounting and connecting the FGB box in the SM (Service Module) in place of its failed counterpart which he has also removed since yesterday. A functions test by TsUP-Moscow over RGS (Russian Groundsite) is scheduled for later today. [The passive KURS-P is required for the Progress 31P docking next week (11/30) at the DC1 nadir port, with KURS backup provided by the manual TORU system.]

Subsequent to the KURS installation and test, Lonchakov has three additional tasks on his schedule:

  • Routine maintenance of the SOZh/ECLSS system, including ASU toilet facilities systems/replaceables. [Regular daily SOZh maintenance consists of replacement of the KTO & KBO solid waste containers, replacement of an EDV-SV waste water and EDV-U urine container, and processing U.S. condensate water as it becomes available in a filled CWC from the Lab humidifier],
  • Daily IMS (Inventory Management System) maintenance, updating/editing its standard “delta file” including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur), and
  • Periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS hatchways, including the passageways PrK (SM Transfer Compartment)–PrK–RO (SM Working Compartment), PkhO (SM Transfer Tunnel)–RO, PkhO–DC1, PkhO–FGB PGO, FGB PGO–FGB GA, FGB GA–Node-1. [This checkup is especially important when the ventilation/circulation system has to cope with a larger crew on board, currently ten persons, and one of the two Russian SKV air conditioners off (SKV-1).]

During the planned reactivation of the Elektron oxygen generator machine at 32 amps by the ground later today (~4:16pm), Lonchakov will monitor the external temperature of its secondary purification unit (BD) for the first 10 minutes of operations to ensure that there was no overheating. [The gas analyzer used on the Elektron during nominal operations for detecting hydrogen (H2) in the O2 line (which could cause overheating) is not included in the control algorithm until 10 minutes after Elektron startup.]

As has become standard operating procedure after deactivation/reactivation of VD-SU monitoring mode, as required for the Elektron power-down/-up, Yuri will also perform a quick function verification of the SUBA Ethernet connection between the OpsLAN (Operations Local Area Network) and the BRI Smart Switch Router in the SM.

CDR Fincke & FE-2-18 Magnus are continuing with WRS (Water Recovery System) activation activities. [The UPA (Urine Processor Assembly) shut down last night after processing pretreated urine for about 2 hrs, probably due to mechanical interference in the DA (Distillation Assembly) centrifuge (one of three). The centrifuge was restarted in maintenance mode and performed nominally. UPA will be restarted today. Meanwhile, the WPA (Water Processing Assembly) continues to collect condensate water and is ready to begin processing when the quantity of distillate collected reaches the minimum amount required to begin.]

Afterwards, Magnus will collect samples of the first WPA-produced water for archival storage and in-flight chemistry/microbiology analysis. Using MCDs (microbial capture devices) from the U.S. WMK (water microbiology kit) for treatment/processing (after no more than 6 hours of the collection), Sandy will then process the samples in the MWAK (microbial water analysis kit) for inflight coliform bacteria (Escherichia coli) detection.

Chamitoff & Magnus have 2:50h set aside for performing Part 1 of installing the PWD (Potable Water Dispenser) hose, today routing the jumper behind several Lab racks which will have to be rotated away from the wall after disengaging their UAMs (Upper Attach Mechanisms).

Other tasks by the FE-2-18 in support of getting the new Regen-ECLSS (Environment Control & Life Support System) on line on today’s schedule include –

  • Completing installation of the Galley (ER6/EXPRESS Rack 6) in the Lab by mating umbilicals and setting up its ELC (ER Laptop Computer for later activation by the ground (POIC/Payload Integration & Operation Center),
  • Processing samples from the successfully installed TOCA for activation & checkout, and
  • Performing leak checks on the WRS-1 & WRS-2 racks.

Gregory & Sandra have another ~4 hrs blocked out for standard joint “handover” activities, to be continued through the docked period ahead.

Mike Fincke is scheduled to feed the butterflies in the CGBA-5/CSI-3 (Commercial Generic Bioprocessing Apparatus/CGBA Science Insert 3) by powering down and decabling the payload, accessing its interior and exposing new food to the larvae in the Butterfly Habitat, then closing up, recabling and reactivating the payload. [The two world-famous spiders meanwhile have torn down their initially confused three-dimensional webs and have produced traditional symmetrical ones, amazing fascinated watchers and puzzling experts with this surprisingly quick adaptation to zero-G. See more below under Science Update.]

The ISS crew completed their physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (FE-2-18), TVIS treadmill (CDR, FE-1, FE-2), RED (CDR, FE-2), and VELO bike with bungee cord load trainer (FE-1). It was the second exercise session for Sandy Magnus, now to become a daily protocol for her.

Cargo Transfers: MPLM (Multi-Purpose Logistics Module) cargo transfers by Pettit, Ferguson, Boe, Kimbrough and Fincke are continuing at a brisk pace, remaining slightly ahead of schedule. [As of this morning, Middeck transfers were 39% complete, MPLM 77%, overall 67%.]

EFBM Contingency Task: During yesterday’s checkout of the Kibo EFBM (External Facility Berthing Mechanism) by Fincke & Magnus, one of the structural latches (SL#1) failed to retract. The IMMT (ISS Mission Management Team) today approved a JAXA request to add a 20-min contingency task to EVA-4 to retract the latch manually.

Shuttle/ISS Reboost: The planned reboost by the Orbiter was performed successfully yesterday at 12:10pm, raising the stack’s mean altitude by 2.2 km.

Weekly Science Update (Expedition Eighteen -- Week 4)

3-D SPACE: Planned.

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Complete.

ANITA (Analyzing Interferometer for Ambient Air): Continuing.

BCAT-3/4 (Binary Colloidal Alloy Test 3/4): No news.

BIO-4: Complete.

CARDIOCOG-2: Complete.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): In progress. For the next CCISS session, Mike will be using a newly refurbished unit that will arrive on ULF-2.

CFE (Capillary Flow Experiment): Reserve.

CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Samples returned on 1J.

CSI-3/CGBA-5 (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus 5): “Mike and Sandy – thank you for the superb job of installing the spiders and larvae in their new environment and for the beautiful HD video of CSI-03 installation. Teachers, students, and all of the CGBA team are enjoying it. The two spiders are cohabitating somewhat peacefully so far and we have witnessed some web-building and tearing down of web, but no direct observation of eating. However, we have observed that the output part of the nutrition cycle is working for at least one of the spiders so there must be ingestion going on at some point. Both have been seen drinking from the water source in the cubbies. The butterfly larvae continue to eat voraciously. We can see only five out of six at any one time, but they tend to crawl everywhere, including inside the food chambers, so it may be that we are looking at a different ‘five’ in each image. Even the fruit flies have supplied some entertainment with their adaptations to micro-g. Students have begun to access the webpage where images are being posted for comparison to the habitats they have in their classrooms. We’ve heard from many teachers about the excitement and enthusiasm these experiments have generated already.”

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): Completed.

EuTEF (European Technology Exposure Facility): On 11/5, EuTEF was finally permanently re-activated, after the PLEGPAY memory erase activity. The MIL-BUS error has occurred 3 times this week. Each time the power feeder#1 is cycled and the instruments are brought back on. There was also a problem on 11/19 with a DMS error at COL-CC which prevented commanding that day.- DEBIE-2: 24-hrs science script ran, started on 11/17 and 11/18, and 11/20. DEBIE-2 is running this week to record any debris impact that may take place in conjunction with the SARJ clean-up activities;- DOSTEL: On-going science acquisition;- EuTEMP: Currently inactive as planned;- EVC: Started picture taking on 11/20;- EXPOSE: On-going science acquisition;- FIPEX: Next script will run after the EVAs are completed;- MEDET: On-going science acquisition; PLEGPAY: Inactive, “Experiment 1” memory has been erased on 10/30. Plasma generation capability has been disabled;- TRIBOLAB: Anomaly with TRIBOLAB Ball Bearing experiment motor, currently being analyzed;

FSL (Fluid Science Laboratory): FSL is nominal.

GEOFLOW: MMA measurements were performed during the docking of ULF2 on 11/17 and successfully downlinked. It is planned to resume the GEOFLOW science runs after 30P undocking, ULF-2 mission and 31P docking.

HDTV System Test DL (JAXA): Complete.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.

Integrated Immune: Underway for Sandra Magnus.

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): In progress.

Marangoni Experiment for ISS in JAXA FPEF (Fluid Physics Experiment Facility): In progress.

Micro-G Clay (JAXA EPO): Complete.

MISSE (Materials ISS Experiment): Ongoing.

Moon Photography from ISS (JAXA EPO): Complete.

MOP (Motion Perception in Zero-G): Three ULF-2 astronauts are participating in this experiment (daily questionnaires)

MSG-SAME (Microgravity Science Glovebox): Complete.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.

MULTIGEN-1: Completed.

MUSCLE-G (LBP/Low Back Pain): Six ULF2 astronauts are participating in this experiment (daily questionnaires)

NOA-1/-2 (Nitric Oxide Analyzer, ESA): Complete.

NUTRITION w/REPOSITORY: “Greg, thanks for the great work in completing all Nutr/Rep inflight sessions. Your participation is greatly appreciated! All samples will be returning with you, allowing the PI to begin analysis soon after return. Mike, Thanks for the great job completing your FD30 session. Crew notes were very helpful.”

PADLES (Passive Dosimeter for Lifescience Experiment in Space): In progress.

PCRF (Protein Crystallization Research Facility) Reconfiguration (JAXA): Complete.

PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SAMPLE: Complete.

SHERE (Shear History Extensional Rheology Experiment): The 25 new Fluid Modules are now on-orbit with STS-126. They are a new-and-improved design that should be easier to deploy and close than the ones Greg Chamitoff used, and researchers are looking forward to performing a new series of experiments with Mike Fincke.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Greg, your only remaining sleep activity will be doffing and stowing the Actiwatch on the Shuttle with the Shuttle crew prior to landing. Mike, thanks for completing the download and initialization activity. Your remaining activities include monthly downloads, 3 weeks of sleep logging and then stowing the Actiwatch. Any additional Sleep logging is above and beyond and greatly appreciated by the PI.”

SOLAR (Solar Monitoring Observatory): SOLAR (USOS) – The Sun visibility window is planned to start on 11/21. Some anomalies are currently being worked for SOLACES and SOVIM instruments. A software patch was successfully uploaded on 11/20 and SOLAR is now running version 4.7 and is in pointing mode;- SOVIM: No science acquisition, non-nominal mode: instrument power consumption was much lower than expected in nominal mode. Trouble-shooting activities were performed on 11/11 and these were not successful. Ground teams are analyzing the data, before next attempts are planned. - SOLSPEC: –Waiting for Sun observation window;- SOLACES: Waiting for Sun observation window.

SOLO (Sodium Loading in Microgravity): Complete.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): Reserve.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: Planned.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): Complete.

CEO (Crew Earth Observations): “There won't be any CEO inputs until 12/1. We stand down from CEO ops and assist with shuttle imagery tasks during shuttle missions.”

No CEO photo targets uplinked for today.

CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 9/1/08, this database contained 770,668 views of the Earth from space, with 324,812 from the ISS alone).

ISS Orbit (as of this morning, 11:13am EST [= epoch]):
Mean altitude -- 353.8 km
Apogee height -- 358.2 km
Perigee height -- 349.4 km
Period -- 91.62 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0006509
Solar Beta Angle -- 19.3 deg (magnitude decreasing)
Orbits per 24-hr. day -- 15.72
Mean altitude gain in the last 24 hours – 2200 m
Revolutions since FGB/Zarya launch (Nov. 98) – 57346.

Significant Events Ahead (all dates Eastern Time, some changes possible.):
11/26/08 -- Progress M-66/31P launch
11/27/08 -- STS-126/Endeavour/ULF-2 undocking; 10:40am – Under Review
11/29/08 -- STS-126/Endeavour/ULF-2 landing; ~2:10pm – Under Review
11/30/08 -- Progress M-66/31P docking – DC1 Nadir (~7:23am)
12/07/08 -- Progress M-65/30P reentry (after 3 weeks autonomous flight for geophysical experiments)
12/18/08 -- Russian EVA-21
02/09/09 -- Progress M-66/31P undocking & deorbit
02/10/09 -- Progress M-67/32P launch
02/12/09 -- Progress M-67/32P docking
02/12/09 -- STS-119/Endeavour/15A launch – S6 truss segment
02/14/09 -- STS-119/Endeavour/15A docking
02/24/09 -- STS-119/Endeavour/15A undocking
02/26/09 -- STS-119/Endeavour/15A landing (nominal)
03/25/09 -- Soyuz TMA-14/18S launch
03/27/09 -- Soyuz TMA-14/18S docking (DC1)
04/05/09 -- Soyuz TMA-13/17S undocking
04/07/09 -- Progress M-67/32P undocking & deorbit
05/15/09 -- STS-127/Endeavour/2J/A launch - JEM EF, ELM-ES, ICC-VLD
05/27/09 -- Soyuz TMA-15/19S launch
Six-person crew on ISS
07/30/09 -- STS-128/Atlantis/17A – MPLM (P), last crew rotation
10/15/09 -- STS-129/Endeavour/ULF3 - ELC1, ELC2
12/10/09 -- STS-130/Endeavour/20A – Node-3 + Cupola
02/11/10 -- STS-131/Atlantis/19A – MPLM(P)
04/08/10 -- STS-132/Endeavour/ULF4 – ICC-VLD, MRM1 (contingency)
05/31/10 -- STS-133/Endeavour/ULF5 – ELC3, ELC4 (contingency).

Note: The daily ISS On-Orbit Status reports can also be found at
http://www.hq.nasa.gov/osf/iss_reports/index.htm

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks